• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 23:50:32 on September27,2006

    Entry number 185625

    keyword=Swing Shift Summary

    Shift Leader: Andrew Puckett
    Target Operator: Michael Paolone
    3rd: Barry Berman

    16:00-17:00: Controlled Access work carrying over from day shift. See JP's entry. Switching DAQ to two-arm coin. trigger. We can't move the right arm to high enough angles to do elastic ep coincidences at the first two left-arm settings, so instead we prescale the right arm T1 and T2 by the maximum amount so that effectively we are running proton singles. However, we still don't prescale T5 (coincidence) or T3 (left arm trigger). Other work performed during the access was to remove the front collimators, change the polarity on the left HRS, change a chiller on the beam calorimeter, and to close the FPP carbon doors.

    17:00-22:00: Run ep elastic at proton angle of 34.38 degrees, proton singles. Target is 15 cm LH2 cell (loop 2). Start with momentum of 755.1 MeV (delta = -2%). Run for ~4 hours with beam current such that T3 rate is about 2.5 kHz.

    from about 20:00 to 20:40 Arne performed a beam charge calibration run consecutively with MCC's PZT test. This period corresponded to CODA run# 5265.

    22:00-00:00: Changed HRSL momentum to 740.3 MeV (delta = 0%). Continued to run on 15 cm LH2 at 2 kHz T3 rate. For the next shift, workers will need to complete the 4 hr. hydrogen run at this momentum, then take short (5 minute) Carbon and dummy cell runs, and then move to hydrogen on the delta=+2% setting before changing kinematics. The right arm is sitting at an unspecified but large (say, 40 +- 10 degrees) angle until we get to the third kinematic of the ep running, at which point we can begin to run ep coincidences. This will probably happen during tomorrow's day shift or swing shift if everything continues to run smoothly, so the angle of the right HRS needs to be known/understood/moved to the right place before then, but for now it is basically just sitting out of the way, and its data acquisition rate is prescaled down to nothing while we take proton singles on the left arm. Xiaodong has informed me that the few beamline HV channels which are "red" and which I was concerned about in my earlier HALOG are unrelated to beam delivery and are not in use/do not concern us at this time.